# Mixtera: A Data Plane for Foundation Model Training

### v2

@@ 6.2. Multimodal LLaVA Finetuning
We are not aware of prior work on the impact of data mixtures on VLMs. To showcase Mixtera’s multimodal capabilities, we evaluate the impact of static mixtures for finetuning a LLaVA-style model (Liu et al., 2024). The LLaVA framework trains an adapter between a pre-trained image encoder and a pre-trained LLM, and then fine-tunes the adapter and LLM on visual instruction-following data.
Training setup. We base our training setup on the TinyLlaVA Factory codebase by Jia et al. (Jia et al., 2024). We rely on a recipe from the Factory and use google/siglip-so400m-patch14-384 (Alabdulmohsin et al., 2023) as the vision encoder, TinyLlama/TinyLlama-1.1B-Chat-v1.0 (Zhang et al., 2024) as the LLM, and a 2-layer MLP as the adapter (Team, 2024a). We train all models on one GH200 node with 4 data parallel GPUs. For pre-training, we use a global batch size of 512 with a learning rate of 0.001, and for finetuning use a global batch size of 128 with a learning rate of 0.00002. We use a chunk size of 256, a cosine learning rate scheduler and the Adam optimizer. We follow Liu et al. (Liu et al., 2024; Team, 2023) and pre-train the adapter on a 558 k subset of the LAION-CC-SBU dataset with BLIP captions. For finetuning, we follow the TinyLLava Factory “LLaVA dataset” and use 665 k samples from six datasets (COCO (Lin et al., 2014), GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and VisualGenome (VG) (Krishna et al., 2017), and LLaVA’s text-only SFT annotations (Liu et al., 2024)) (Team, 2024b). We pre-train the adapter once, and then vary the proportions of the finetuning datasets. We randomly generate 256 mixtures for finetuning. Since the number of datapoints in comparison to LLM training is small, we use a best-effort mixture and ensure we go through all samples exactly once. All models see the same data, but in different order.
@@
Results. In Table 4, we show the results reported by Jia et al. (Jia et al., 2024), the results we obtain using the inferring mixture (Section 5.2.2), and three mixtures out of the generated mixtures that perform well. While the inferring mixture does not achieve their reported results, this could be either due to different data dynamics in their training, or due to a different evaluation setup we cannot reproduce as their model weights are not public. Nevertheless, in particular on MME/MMMU/POPE, the mixtures outperform the baseline. Mix. #252 is weighted towards TextVQA (29.1 %) and OCR-VQA (19.5 %), with equal proportions of COCO (21.9%) and VG (21.9 %); GQA (4.0 %) and SFT annotations (3.6 %) contribute minimally. Mix. #107 places a large emphasis on SFT annotations (28.9 %) and VG (27.4 %), followed by COCO (25.5 %); GQA (5.3 %) and TextVQA (5.1 %) have lower representation, while OCR-VQA (7.8 %) remains a minor component. Mix. #155 prioritizes OCR-VQA (30.3 %) and TextVQA (25.7 %), and SFT annotations (22.7 %); VG (13.2 %) and COCO (6.9 %) are underrepresented, while GQA (1.2 %) is least utilized. Overall, despite seeing the same samples globally, the mixtures play a big role for model accuracy. Exploring mixtures for pre-training of VLMs and adapting dynamic mixtures to VLMs are promising future work directions.
@@ 6.3. Throughput Benchmarks
We now focus on throughput and compare Mixtera with other data loaders across various configurations. Mixtera’s goal is to avoid data stalls, i.e., training throughput should not be reduced because we are waiting for data (Mohan et al., 2021; Murray et al., 2021; Kuchnik et al., 2022; Zhao et al., 2022). Hence, we measure throughput in tokens per second in actual workloads. Note that smaller models and more data parallelism increase the pressure on the data loader, while larger models reduce the pressure. If a data loader can sustain training small models at scale on a high-end platform like the GH200, its performance is sufficient for other scenarios as well.
@@
Besides Mixtera with chunk sizes of 512, 1024, and 2048, we benchmark the iterable HuggingFaceDataset by Torchtitan (Hf-Iter), a mapped version (Hf-Map), the Mosaic StreamingDataset (Team, 2022) (Mosaic). The difference between Hf-Iter and Hf-Map is that similar to Mixtera, Hf-Iter loads and tokenizes the data on the fly, while Hf-Map preprocesses all data, including tokenization. We evaluate throughput on the 162M model since larger models only lead to lower throughput. We always use FSDP since Torchtitan only enables bfloat16 training if sharding is used. We activate compilation, disable activation checkpointing, and use fused AdamW. We measure throughput for 30 steps, discarding the first step. We repeat all measurements three times, i.e., in total we have 3x30 steps. We use the huggingface EleutherAI/gpt-neox-20b tokenizer for all data loaders. We store all data on the Iopsstor SSD-backed Lustre DFS in the Alps cluster.
TODO!
@@
Single-node. We train on a single GH200 node with 4 GPUs. We use 2 data parallel replicates and shards. In Figure 7 we show the throughput for the data loaders depending on the number of data workers. We test up to 16 workers, where 0 workers indicate that data is loaded in the same process as the main training loop. We only show results up to 4 workers due as throughput does not increase further. The error bars show the standard deviation of the throughput. Without data workers, Mixtera has the highest average throughput; with one worker, the other data loaders reach their peak performance.
@@
Mixtera performs worse than the other loaders with a lower number of workers due to the random access into the files. For every sample, it needs to open the file, seek to the correct position, and load the data, instead of bulk-transferring all the data as the other data loaders can do. This leads to the higher variance in throughput indicated by the error bars, which overall leads to a lower average. However, with a higher number of workers, the variance in data access gets hidden. When using 4+ workers, Mixtera performs almost identical to the other data loaders. Additionally, increasing the chunk size also helps, and for the 0 worker case, Mixtera with a chunk size of 1024 or 2048 even has a higher average throughput.
This benchmark setup is quite extreme, as we train a very small 162M model on an extremely fast GPU. Nevertheless, just by adding some workers, which comes at no overhead beyond some CPU resources, Mixtera reaches competitive throughput despite its more complicated model.
@@
Scaling out. We investigate how the data loaders scale for larger training jobs with higher data parallelism. We scale up to 64 GH200 nodes with a total of 256 data parallel GPUs. We find that using a maximum number of 16 replication GPUs works best. For 4, 8, and 16 GPUs, we use half the GPUs as replication, and shard across the rest. The results with 8 data workers can be found in Figure 8. All data loaders scale perfectly linear. For larger number of GPUs, the throughput variance increases a bit for all systems. We attribute this to the random assignment of nodes in the cluster by the slurm scheduler across the 3 repetitions. We do not test pipeline or tensor parallelism as (i) this is not necessary for the 162M model (and a larger model would only stress the data loaders less), (ii) the data loaders besides Mixtera do not easily support 3D parallelism, (iii) increasing data parallelism increases the load on the system more.
@@
TODO!
6.3.1. Query execution
Mixtera executes the query in DuckDB and generates the ChunkerIndex before we can start training. On our Pile dataset with 241 M samples, DuckDB takes 29 s, and preparing the index takes 18 s due to our C++ implementation. Optionally, we can also persist an initial state checkpoint of the query, which due to serialization of nested Python dictionaries takes 13 min. This is only necessary if checkpointing should be supported. Future checkpoints can then be stored in milliseconds. While the streaming Hf-Iter data loader can basically start streaming data immediately, the Hf-Map data loader, the default in nanotron, loads and tokenizes the data first, taking 2 h 51 min for this dataset. However, what needs to be considered is that for streaming data loaders such as Hf-Iter or Mosaic, in many scenarios, users would also need to run more offline preprocessing, e.g., to perform static filtering, or reshuffling the data if we want to mix on a different property. Mixtera avoids this offline preprocessing completely, and, besides an optional state checkpoint, starts streaming data in under a minute.
@@ 7. Conclusion and Future Work
Managing training data at scale requires data management support. We present Mixtera, a training framework-agnostic data plane for foundation model training. Using its chunking mechanism, it supports dynamic mixtures across arbitrary properties. We demonstrate Mixtera’s performance and scalability, as well as the importance of mixtures on both LLMs and VLMs. For future work, it is interesting to extend the analysis to the cloud, i.e., using a service like S3 as the underlying storage layer. Mixtera also lays the foundation to implement lineage tracking on which model was trained on which data, as it is a centralized access point. From a ML perspective, the impact of data composition and quality on model performance is still under-explored, and Mixtera provides a foundation to conduct future research on dynamic and static mixing algorithms.
@@ Acknowledgements.
We thank Beste Güney for her contributions to Mixtera’s codebase. We thank Antoni-Joan Solergibert i Llaquet, Imanol Schlag, Steven Hand, Martin Jaggi, Antoine Bosselut, Alexander Hägele, Loubna Ben Allal, Quentin de Laroussilhe, Paul Barham, Yiding Jiang, Theodoros Rekatsinas, and Foteini Strati for helpful discussions. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a09 on Alps. Maximilian Böther is supported by the Swiss National Science Foundation (project number 200021_204620).References
@@
- Armbrust et al. (2021) Michael Armbrust, Ali Ghodsi, Reynold Xin, and Matei Zaharia. 2021. Lakehouse: A New Generation of Open Platforms that Unify Data Warehousing and Advanced Analytics. In Proceedings of the Conference on Innovative Data Systems Research (CIDR).
@@
- ETH AI Center & EPFL AI Center (2025) ETH AI Center & EPFL AI Center. 2025. Swiss AI Initiative. https://www.swiss-ai.org/.
@@
- Fusco et al. (2024) Luigi Fusco, Mikhail Khalilov, Marcin Chrapek, Giridhar Chukkapalli, Thomas C. Schulthess, and Torsten Hoefler. 2024. Understanding Data Movement in Tightly Coupled Heterogeneous Systems: A Case Study with the Grace Hopper Superchip. arXiv preprint (2024). doi:10.48550/ARXIV.2408.11556
@@
- Liu et al. (2023) Hanmeng Liu, Jian Liu, Leyang Cui, Zhiyang Teng, Nan Duan, Ming Zhou, and Yue Zhang. 2023. LogiQA 2.0—An Improved Dataset for Logical Reasoning in Natural Language Understanding. IEEE/ACM Transactions on Audio, Speech, and Language Processing 31 (2023). doi:10.1109/taslp.2023.3293046

### v3

@@ 6.2. Multimodal LLaVA Finetuning
To showcase Mixtera’s multimodal capabilities, we evaluate the impact of static mixtures for finetuning a LLaVA-style model (Liu et al., 2024). We are not aware of prior work on the impact of data mixtures on VLMs. The LLaVA framework trains an adapter between a pre-trained image encoder and a pre-trained LLM, and then fine-tunes the adapter and LLM on visual instruction-following data.
Training setup. We base our training setup on the TinyLlaVA Factory codebase by Jia et al. (Jia et al., 2024). We rely on a recipe from the Factory and use google/siglip-so400m-patch14-384 (Alabdulmohsin et al., 2023) as the vision encoder, TinyLlama/TinyLlama-1.1B-Chat-v1.0 (Zhang et al., 2024) as the LLM, and a 2-layer MLP as the adapter (Team, 2024a). We train all models on one GH200 node with 4 data parallel GPUs. For pre-training, we use a global batch size of 512 with a learning rate of 0.001, and for finetuning use a global batch size of 128 with a learning rate of 0.00002. We use a chunk size of 256, a cosine learning rate scheduler and the Adam optimizer. We follow Liu et al. (Liu et al., 2024; Team, 2023) and pre-train the adapter on a 558 k subset of the LAION-CC-SBU dataset with BLIP captions. For finetuning, we follow the TinyLLava Factory (Team, 2024b) “LLaVA dataset” and use 665 k samples from six datasets (COCO (Lin et al., 2014), GQA (Hudson and Manning, 2019), OCR-VQA (Mishra et al., 2019), TextVQA (Singh et al., 2019), and VisualGenome (VG) (Krishna et al., 2017), and LLaVA’s text-only SFT annotations (Liu et al., 2024)). We pre-train the adapter once, and then vary the proportions of the finetuning datasets. We randomly generate 256 mixtures for finetuning. Since the number of datapoints in comparison to LLM training is small, we use a best-effort mixture and ensure we go through all samples exactly once. All models see the same data, but in a different order.
@@
Results. In Table˜4, we show the results reported by Jia et al. (Jia et al., 2024), the results we obtain using the inferring mixture (Section˜5.2.2), and three mixtures out of the generated mixtures that perform well. While the inferring mixture does not achieve their reported results, this could be either due to different data dynamics in their training, or due to a different evaluation setup we cannot reproduce as their model weights are not public. Nevertheless, in particular on MME/MMMU/POPE, the mixtures outperform the baseline. Mix. #252 is weighted towards TextVQA (29.1 %) and OCR-VQA (19.5 %), with equal proportions of COCO (21.9%) and VG (21.9 %); GQA (4.0 %) and SFT annotations (3.6 %) contribute minimally. Mix. #107 places a large emphasis on SFT annotations (28.9 %) and VG (27.4 %), followed by COCO (25.5 %); GQA (5.3 %) and TextVQA (5.1 %) have lower representation, while OCR-VQA (7.8 %) remains a minor component. Mix. #155 prioritizes OCR-VQA (30.3 %) and TextVQA (25.7 %), and SFT annotations (22.7 %); VG (13.2 %) and COCO (6.9 %) are underrepresented, while GQA (1.2 %) is least utilized. Overall, despite seeing the same samples globally, the mixtures play a big role for model accuracy.
@@ 6.3. Throughput Benchmarks
We now focus on throughput and compare Mixtera with other data loaders across various configurations. The goal is avoiding data stalls, i.e., training throughput should not be reduced because the GPU is waiting for data (Mohan et al., 2021; Kuchnik et al., 2022; Murray et al., 2021; Zhao et al., 2022; Graur et al., 2022; Böther et al., 2025; Robroek et al., 2026; Chen et al., 2025). We want to show that Mixtera enables on-the-fly data streaming with dynamic mixing without lowering throughput in comparison to other data loaders. We measure throughput in tokens per second. Note that smaller models and more data parallelism increase the pressure on the data loader, while larger models reduce the pressure as the training computation takes longer. If a data loader can sustain training small models at scale on a high-end platform like the GH200, its performance is sufficient for other scenarios as well.
@@
We run Mixtera with chunk sizes of 512, 1024, and 2048. We compare to well-known state-of-the-art data loaders, i.e., the iterable HuggingFaceDataset by Torchtitan (Hf-Iter), a mapped version (Hf-Map), and the Mosaic StreamingDataset (Team, 2022) (Mosaic). The difference between Hf-Iter and Hf-Map is that similar to Mixtera, Hf-Iter loads and tokenizes the data on the fly, while Hf-Map preprocesses all data, including tokenization. Note that none of these data loaders support dynamic mixing (Table˜1), they just read files from front to back. We evaluate throughput on the 162M model since larger models only lead to lower throughput, as discussed above. We always use FSDP since Torchtitan only enables bfloat16 training if sharding is used. We activate compilation, disable activation checkpointing, and use fused AdamW. We measure throughput for 30 steps, discarding the first step. We repeat all measurements three times, i.e., in total we have 3x30 steps. We use the Hugging Face EleutherAI/gpt-neox-20b tokenizer for all data loaders. We store all data on an SSD-backed Lustre DFS.
@@
Single-node. We train on a single GH200 node with 4 GPUs. We use 2 data parallel replicates and shards. In Figure˜7 we show the throughput for the data loaders depending on the number of data workers, i.e., data loading processes from the torch.DataLoader. We test up to 16 workers, where 0 workers indicate that data is loaded in the same process as the main training loop. The plot only show results up to 4 workers as throughput does not increase further. Without data workers, Mixtera has the highest average throughput; with one worker, the other data loaders reach their peak performance.
@@
Overall, Mixtera provides similar throughput to the baselines while having a much richer feature set, e.g., dynamic mixing. Notably, Mixtera has higher throughput variance than other loaders for lower number of workers. This is due to the random access into the files. For every sample, it needs to open the file, seek to the correct position, and load the data, instead of bulk-transferring all the data as the other data loaders can do. This leads to the higher variance in throughput indicated by the error bars, which overall leads to slightly lower averages. With a higher number of workers and larger chunk sizes, the variance decreases. When using 4+ workers, Mixtera’s throughput matches the throughput of the baselines. Additionally, increasing the chunk size also helps, and for the 0 worker case, Mixtera with a chunk size of 1024 or 2048 even has a higher average throughput.
This benchmark setup is quite extreme, as we train a very small 162M model on an extremely fast GPU. For larger, state-of-the-art model sizes, the throughput differences disappear, as computation time spent within the model forward and backward passes increases. Mixtera’s client-side overhead is minimal; the primary operations (chunk parsing and data fetching) introduce negligible overhead.
@@
Scaling out. We investigate how the data loaders scale for larger training jobs with higher data parallelism. We scale up to 64 GH200 nodes with a total of 256 data parallel GPUs. We find that using a maximum number of 16 replication GPUs works best. For 4, 8, and 16 GPUs, we use half the GPUs as replication, and shard across the rest. Figure˜8 shows the results with 8 data workers. All data loaders scale linearly. With more GPUs, the throughput variance increases a bit for all systems. We attribute this to the random assignment of nodes in the cluster by the Slurm scheduler across the 3 repetitions. We do not test pipeline or tensor parallelism as (i) this is not necessary for the 162M model (and a larger model would only stress the data loaders less), (ii) the data loaders besides Mixtera do not easily support 3D parallelism, (iii) increasing data parallelism increases the load on the system more. This demonstrates the scalability of Mixtera’s single-controller design and implementation, as well as the efficiency of its chunk generation, enabling it to effectively supply chunks to all clients even at scale.
@@
Predicates. The previous benchmarks select all domains within The Pile. To show how a filtering predicate impacts throughput, in Figure˜9, we show the throughput for chunks sizes 512/2048, and 0/2 data workers, across four representative predicates: Books3 only (large samples), CommonCrawl (CC) and ArXiv (frequent domains with small and large samples), CC only (frequent but small samples), and a set of infrequent domains (Enron Emails, NIH ExPorter, PhilPapers, EuroParl, USPTO).
With 0 data workers, throughput varies with the predicate; more workers stabilize throughput. For chunk size 512, Only CC achieves ~587 kTokens/s, while Books3 reaches ~776 kTokens/s. These differences stem from different file access patterns. Because The Pile is shuffled, large domains are more likely to appear in clusters, reducing repeated file scans. Domains also differ in sample sizes: large text samples yield multiple training examples from a single read, lowering total data transfer, as seen with Books3. With two workers, Mixtera masks these effects and throughput is uniform.
Clustering. If data is clustered, Mixtera can generate longer intervals and thereby reduce the amount of data transfers. Figure˜10 compares the throughput of the regular benchmarking data with a clustered version that sequentially writes out each domain’s data. For a small number of workers, we generally observe an increase in avg. throughput and a slight decrease of variance. As before, using more workers hides these effects.
6.4. Data Ingestion
Before executing queries, the user must ingest metadata (c.f. Section˜5.1). During ingestion, Mixtera reads all samples, extracts their metadata, and inserts it into the underlying metadata database. This cost is paid only once; thereafter, arbitrarily many queries can be executed. In Figure˜11 we report time spent reading the samples (scanning all files) and inserting metadata. We benchmark this for varying dataset sizes (10 %, 100 %, and 250 % of The Pile) and number of attributes (randomly generated to avoid correlation). Other operations (e.g., parser setup) have negligible overhead.
The timings are obtained using Mixtera’s optimized default settings: multithreaded file scanning (MT) and an insertion chunk size of 2,000. Disabling MT increases the ingestion time from 315 to around 3,000 seconds for the standard 100 % / 1 attribute case. With MT enabled, decreasing the chunk size to 500 increases ingestion time from 315 s to 340 s. Chunk sizes larger than 2,000 yield only marginal benefit and increase the risk of out-of-memory errors.
6.5. Query Execution Latency Breakdown
When a query is executed in Mixtera, the main steps are querying the populated ODC (DuckDB), building the ChunkerIndex, and, optionally, writing the first checkpoint. In Figure˜12, we give a time breakdown for varying dataset sizes (10 %, 100 %, and 250 % of The Pile) and number of properties (randomly generated to avoid correlation), following Section˜6.4. For the 100 % / 1 attribute case, DuckDB takes 32 s, and preparing the index takes 16 s due to our C++ implementation. Persisting the initial state checkpoint of the query dominates the runtime with 282 s. This is because we have to serialize nested Python dictionaries, which is slow despite engineering optimizations. Importantly, after this initial checkpoint, future checkpoints at the server can be stored within milliseconds, since we do not have to serialize the index again. If no Mixtera checkpoints are needed, the serialization can be skipped.
Other systems. The streaming Hf-Iter data loader can basically start streaming data immediately, the Hf-Map data loader, the default in nanotron, loads and tokenizes the data first, taking 2 h 51 min for The Pile 100 %. However, to use streaming data loaders such as Hf-Iter or Mosaic, in many scenarios, users would also need to run more offline preprocessing, e.g., to perform static filtering, or reshuffling the data if we want to mix on a different property. Mixtera avoids this offline preprocessing completely.
@@ 7. Conclusion and Future Work
Understanding how data mixture recipes affect model quality is an active and critical area of ML research. We design Mixtera to enable researchers and model developers to easily and quickly train models with a variety of mixtures across arbitrary data properties and dynamically vary mixtures during training. Mixtera is a declarative data plane for foundation model training that is training framework-agnostic. We demonstrate Mixtera’s throughput and scalability, as well as the impact of mixtures on model quality for both LLMs and VLMs. The system lays the foundation for implementing additional features, such as data lineage tracking for model training, as it has a global view of the data, and enables future research on compute-intensive modalities such as video (Ye et al., 2025) and audio (Zhao et al., 2025).
@@ Acknowledgements.
We thank Beste Güney for her contributions to Mixtera’s codebase. We thank Antoni-Joan Solergibert i Llaquet, Imanol Schlag, Steven Hand, Martin Jaggi, Antoine Bosselut, Alexander Hägele, Loubna Ben Allal, Quentin de Laroussilhe, Paul Barham, Yiding Jiang, Theodoros Rekatsinas, Gustavo Alonso, and Foteini Strati for helpful discussions. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a09 on Alps. Maximilian Böther is supported by the Swiss National Science Foundation (project number 200021_204620).References
@@
- Barham et al. (2022) Paul Barham, Aakanksha Chowdhery, Jeff Dean, Sanjay Ghemawat, Steven Hand, Dan Hurt, Michael Isard, Hyeontaek Lim, Ruoming Pang, Sudip Roy, Brennan Saeta, Parker Schuh, Ryan Sepassi, Laurent El Shafey, Chandramohan A. Thekkath, and Yonghui Wu. 2022. Pathways: Asynchronous Distributed Dataflow for ML. In Proceedings of the Conference on Machine Learning and Systems (MLSys).
@@
- Bodensohn et al. (2025) Jan-Micha Bodensohn, Ulf Brackmann, Liane Vogel, Anupam Sanghi, and Carsten Binnig. 2025. Unveiling Challenges for LLMs in Enterprise Data Engineering. arXiv preprint (2025). doi:10.48550/arXiv.2504.10950
@@
- Böther et al. (2025) Maximilian Böther, Ties Robroek, Viktor Gsteiger, Xianzhe Ma, Pınar Tözün, and Ana Klimovic. 2025. Modyn: Data-Centric Machine Learning Pipeline Orchestration. In Proceedings of the Conference on Management of Data (SIGMOD). doi:10.1145/3709705
@@
- Chen et al. (2025) Zihao Chen, Chenyang Zhang, Chen Xu, Zhao Zhang, Jiaqiang Wang, Weining Qian, and Aoying Zhou. 2025. Scheduling Data Processing Pipelines for Incremental Training on MLP-based Recommendation Models. In Companion of the International Conference on Management of Data (SIGMOD). doi:10.1145/3722212.3724454
@@
- Graur et al. (2022) Dan Graur, Damien Aymon, Dan Kluser, Tanguy Albrici, Chandramohan A Thekkath, and Ana Klimovic. 2022. Cachew: Machine Learning Input Data Processing as a Service. In Proceedings of the USENIX Annual Technical Conference (ATC).
@@
- Robroek et al. (2026) Ties Robroek, Neil Kim Nielsen, and Pınar Tözün. 2026. TensorSocket: Shared Data Loading for Deep Learning Training. In Proceedings of the Conference on Management of Data (SIGMOD). doi:10.1145/3749185
@@
- Ye et al. (2025) Juncheol Ye, Seungkook Lee, Hwijoon Lim, Jihyuk Lee, Uitaek Hong, Youngjin Kwon, and Dongsu Han. 2025. SAND: A New Programming Abstraction for Video-based Deep Learning. In Proceedings of the Symposium on Operating Systems Principles (SOSP). doi:10.1145/3731569.3764847
@@
- Zhao et al. (2025) Juntao Zhao, Qi Lu, Wei Jia, Borui Wan, Lei Zuo, Junda Feng, Jianyu Jiang, Yangrui Chen, Shuaishuai Cao, Jialing He, Kaihua Jiang, Yuanzhe Hu, Shibiao Nong, Yanghua Peng, Haibin Lin, Xin Liu, and Chuan Wu. 2025. OVERLORD: Ultimate Scaling of DataLoader for Multi-Source Large Foundation Model Training. arXiv preprint (2025). doi:10.48550/arXiv.2504.09844